# The Leftmost particle of branching subordinators

### v1

@@ 1.1.2 $\alpha$-stable trajectories with rare infinite branchings
More precisely, for $\alpha\in(0,1)$, we define by $\nu_{\alpha}$ the normalized $\alpha$-stable Lévy measure on $(0,\infty)$, that is $\nu_{\alpha}(\mathrm{d}\mathfrak{z}):=\frac{\alpha}{\Gamma(1-\alpha)}\frac{\mathrm{d}\mathfrak{z}}{\mathfrak{z}^{\alpha+1}}$. We denote by $\tilde{\nu}_{\alpha}$ the measure on $\mathscr{Q}$ defined on cylinders as follows: for any $k\geq 2$ and $A_{1},\cdots,A_{k}\in(0,\infty],$
$\tilde{\nu}_{\alpha}(A_{1}\times\cdots\times A_{k}\times(0,\infty]\times\cdots):=\nu_{\alpha}(A_{1})\prod_{i=2}^{k}\mathds{1}_{\{\infty\in A_{i}\}},$
[…]
@@ 1.1.3 $\alpha$-stable trajectories with many binary branchings
More precisely, for $\alpha\in(0,1)$ we let $\tilde{\nu}_{\alpha}$ be as in Section 1.1.2. For $\theta>1$, we let $\mu_{\theta}$ be the sigma-finite measure on $\mathbb{R}_{+}$ defined by $\mu_{\theta}(\mathrm{d}\mathfrak{z}):=e^{\mathfrak{z}^{\theta}}\mathds{1}_{\{\mathfrak{z}\geq 0\}}\mathrm{d}\mathfrak{z}$. We denote by $\tilde{\mu}_{\theta}$ the measure on $\mathscr{Q}$ defined on cylinders as follows: for any $k\geq 3$ and $A_{1},\cdots,A_{k}\in(0,\infty]$,
$\tilde{\mu}_{\theta}(A_{1}\times\cdots\times A_{k}\times(0,\infty]\times\cdots):=\mathds{1}_{\{0\in A_{1}\}}\mu_{\theta}(A_{2})\prod_{i=3}^{k}\mathds{1}_{\{\infty\in A_{i}\}},$
[…]
The measure $\Lambda^{mb}_{\alpha,\theta}$ clearly falls in the case (8) so this example illustrates that, for branching subordinators, one can very well have only binary branchings and yet be in the case (8), which would not be possible for branching random walks. This shows that the setting of branching subordinators is somehow more adapted than the setting of branching random walks to generate natural examples in the case (8) which, as we will see in Section 1.3, allows for a larger variety of asymptotic behaviours, compared with the classical case (6).
@@ 1.2 Some more definitions
The function $\kappa_{a}(\cdot)$, $a>0$, has a natural interpretation. For any $\bm{x}=(x_{i})_{i\geq 1}\in\mathscr{Q}$, we set $T_{a}:\bm{x}\in\mathscr{Q}\mapsto\bm{x}^{a}\in\mathscr{Q}$ to be the $a$-truncated operator on $\mathscr{Q}$, where $\bm{x}^{a}$ is defined to be the sequence $\bm{x}$ capped at $a$ such that $x_{1}$ remains unchanged, that is $(\bm{x}^{a})_{1}:=x_{1}$ and for any $i\geq 2$, $(\bm{x}^{a})_{i}:=x_{i}\mathds{1}_{\{x_{i}<a\}}+\infty\mathds{1}_{\{x_{i}\geq a\}}$. Let $\Lambda^{a}(\mathrm{d}\bm{x}):=\mathds{1}_{\{\bm{x}\neq(0,\emptyset)\}}(T_{a}\Lambda)(\mathrm{d}\bm{x})$, where $T_{a}\Lambda$ denotes the image measure of $\Lambda$ by $T_{a}$. Since $\lim_{a\to\infty}\Lambda^{a}(\mathscr{Q}^{\star})>0$ by (3), there exists $a_{0}(\Lambda)\geq 0$ such that $\Lambda^{a}(\mathscr{Q}^{\star})>0$ for all $a>a_{0}(\Lambda)$. Note that if $\Lambda$ satisfies (3) and (4), then $\Lambda^{a}$ satisfies the conditions (3) and (6) for all $a>a_{0}(\Lambda)$. For a branching subordinator $Y$ with characteristics $(d,\Lambda)$ and $a>a_{0}(\Lambda)$, we denote by $Y^{a}$ the sub-system of $Y$ obtained as follows: at each branching event (represented by an $\bm{x}=(x_{i})_{i\geq 1}\in\mathscr{Q}$ with $x_{2}<\infty$), the particle born with displacement $x_{1}$ is kept and for $i\geq 2$, the particles born with displacement $x_{i}\geq a$ are removed from the system, along with their future lineage. Then $Y^{a}$ is a branching subordinator with characteristics $(d,\Lambda^{a})$ and the corresponding Laplace exponent is $\kappa_{a}(\cdot)$, that is, for any $t,\lambda\geq 0$
$\mathbb{E}\Big{[}\sum_{u\in\mathscr{N}^{a}(t)}e^{-\lambda Y^{a}_{u}(t)}\Big{]}=e^{-t\kappa_{a}(\lambda)}\in(0,\infty),$
where, similarly as for $Y$, $\mathscr{N}^{a}(t)$ stands for the set of particles in $Y^{a}$ alive at time $t$ and for any particle $u\in\mathscr{N}^{a}(t)$ and $s\leq t$, $Y^{a}_{u}(s)$ denotes the position of $u$ at time $s$, if already born at time $s$. Otherwise, $Y^{a}_{u}(s)$ stands for the position of the most recent ancestor of $u$ alive at time $s$. We refer to Appendix A for a joint construction of $Y$ and $Y^{a}$ where $Y^{a}$ is introduced first in Definition A.1, $Y$ is subsequently constructed in Definition A.4, and a slightly different version of the above restriction property is stated in Lemma A.3. Also note that each canonical trajectory associated with $Y^{a}$ is a subordinator with Laplace exponent $\phi$, see below (3) for a description of canonical trajectories and Definition A.2 for a rigorous definition.
@@ 1.3 Main results
The main purpose of the present paper is to study the behaviour of the leftmost particle of branching subordinators. This behaviour is of particular interest to us because it turns out that, under our assumptions, for a branching subordinator $Y$ with characteristics $(d,\Lambda)$, the growth of $t\mapsto\underline{Y}(t)=\min_{u\in\mathscr{N}(t)}Y_{u}(t)$ can possibly display interesting sub-linear behaviors.
Introduce $\underline{Y}(\infty):=\lim_{t\to\infty}\underline{Y}(t)$. Note that $\underline{Y}(\infty)$ is well defined since the function $t\mapsto\underline{Y}(t)$ is non-decreasing. We start by giving a criteria in terms of the truncated Laplace exponent to determine whether $\underline{Y}(\infty)$ is finite or infinite. For this, we first note from (4) and dominated convergence that, for any $0<a_{1}<a_{2}$, $M_{a_{2}}(\lambda)-M_{a_{1}}(\lambda)$ converges to 00 as $\lambda$ goes to infinity, so the quantity
@@
is independent of the choice of $a>0$.
@@
If the assumption (6) holds true for some $\theta\geq 0$ (we recall that in this case the Laplace exponent $\kappa(\cdot)$ introduced in (7) is well-defined on $[\theta,\infty)$), then we have $\mathscr{L}(d,\Lambda)=\lim_{\lambda\rightarrow\infty}\kappa(\lambda)$.
Our criteria for finiteness or infiniteness of $\underline{Y}(\infty)$ is as follows.
@@
The event $\{\underline{Y}(\infty)<\infty\}$ is referred to as the phenomenon of explosion in [1]. Focused on branching random walks with non-negative displacements, authors provide conditions for explosion to happen. In our case, in view of Proposition 1.3, it is sufficient to assume $\mathscr{L}(d,\Lambda)>0$ to ensure that explosion does not occur. A natural question is that of the rate of growth of $\underline{Y}(t)$. In any case, under (3) and (4), we have, $\mathbb{P}$-almost surely
@@
As for branching random walks (see for example the famous works of Hammersley [16], Kingman [18] and Biggins [10]), if (6) holds for some $\theta\geq 0$ and if we further require that $-\infty<\kappa(\theta_{0})<0$ for some $\theta_{0}>0$ (allowing $\kappa(0)=-\infty$), then the limit exists and is nonrandom, more precisely, we have $\mathbb{P}$-almost surely
$\lim_{n\to\infty}\frac{\underline{Y}(t)}{t}=\sup_{\lambda\geq\theta_{0}}\frac{\kappa(\lambda)}{\lambda}=\sup_{\lambda>0}\frac{\kappa(\lambda)}{\lambda}\in[0,\infty).$ (14)
Therefore, in the latter case and in view of (14), the growth of $t\mapsto\underline{Y}(t)$ is linear if and only if $\mathscr{L}(d,\lambda)>0$; in particular, in view of Proposition 1.3 and excluding the special case $\mathscr{L}(d,\Lambda)=0$, we have that, either $\underline{Y}(t)$ is of order $t$ as $t\to\infty$ or $\underline{Y}(\infty)\in[0,\infty)$.
Note that, by (13), the classical linear growth is observed when $d>0$, even in the case (8), that is when the exponential moments assumption (6) does not hold. We discuss the linear behavior and provide a brief justification of (13) and (14) in Appendix B.
@@
We define a measure $\mu$ on $[0,\infty)$ by
$\forall A\in\mathscr{B}(\mathbb{R}_{+}),\ \mu(A):=\int_{\mathscr{Q}}\left(\sum_{k=2}^{\infty}\mathds{1}_{\{x_{k}\in A\}}\right)\Lambda(dx).$ (16)
The measure $\mu$ can be understood as a measure-valued average of the contributions of branchings. We note that $M_{a}(\lambda)=\int_{[0,a)}e^{-\lambda\mathfrak{z}}\mu(\mathrm{d}\mathfrak{z})$. If we had $\mu([0,a))$ of order $e^{ca}$, then (6) would be satisfied for any $\theta>c$. Since we are mostly interested in the case (8), we assume an higher rate of growth for $\mu([0,a))$, but also some regularity. More precisely, our second assumption is the following.
@@
$M_{2,a}(\lambda):=\int_{\mathscr{Q}}\left(\sum_{k=2}^{\infty}e^{-\lambda x_{k}}\mathds{1}_{\{x_{k}<a\}}\right)^{2}\Lambda(dx).$ (18)
@@
Remark 1.4 (Binary/trinary branchings).
If the system has only binary branchings, that is, if $\Lambda(\{\bm{x}\in\mathscr{Q},x_{3}<\infty\})=0$, then we have $M_{2,a}(\lambda)=M_{a}(2\lambda)$ for any $a,\lambda\geq 0$. If the system has only binary and trinary branchings, that is, if $\Lambda(\{\bm{x}\in\mathscr{Q},x_{4}<\infty\})=0$, then one can easily see that we have $M_{2,a}(\lambda)\leq 2M_{a}(2\lambda)$ for any $a,\lambda\geq 0$. This will allow to conveniently check that Assumption 3 is satisfied in some examples involving binary or trinary branchings (see Section 2.2 and Remark 2.2).
@@
Let $\Lambda$ be a measure on $\mathscr{Q}$ satisfying (3)-(5) and Assumption 1 for some $\alpha\in(0,1)$, Assumption 2 for some $\sigma>0$, and Assumption 3. Then the branching subordinator with characteristic couple $(0,\Lambda)$ satisfies (15) with $\gamma:=\frac{1}{\alpha+(1-\alpha)(1+\sigma)}$.
@@
We will see in Remark 2.2 that the result of Corollary 1.7 remains true with the same value of $\gamma$ when the binary branchings from the example of Section 1.1.3 are replaced by trinary branchings.
The fact that the growth of $t\mapsto\underline{Y}(t)$ is slower than linear under our assumptions can informally be explained as follows: in cases where $\lim_{a\to\infty}M_{a}(\lambda)=\infty$ for any $\lambda\geq 0$, each particle in $Y$ is somehow allowed to give progeny to a very large number of children (see Example 1.1.2 for instance) or/and to create newborn children very often (see Example 1.1.3 for instance), producing a population of particles sufficiently large to maintain $\underline{Y}(t)$ below $\mathfrak{c}t$ for any $\mathfrak{c}>0$. Besides, one can notice that the rate of growth of $t\mapsto\underline{Y}(t)$ strongly depends on $\alpha$ (see Assumption 1) and $\sigma$ (see Assumption 2). The parameter $\alpha$ acts on the intensity of the (infinitely many) small jumps of each particle whereas $\sigma$ acts on the distribution of the positions of the children of each particle and the combination of both is largely responsible for this rate of growth.
The idea of the proof of Theorem 1.5 is as follows. Proving the upper bound relies on estimating the average number of slow particles. Proving the lower bound is more involved and relies on understanding the mechanism by which slow particles are generated and to show that, in some sense, the structure of branchings allows for the set of slow particles to sustain itself over time. In other words we produce events of large probability on which slow trajectories produce sufficiently enough slow particles. This requires to prove a large deviation estimate that give a lower bound on the probability for a particle to be slow, and to show that there is some regularity on the distribution of particles that are born from the slow trajectories.

### v2

@@ 1.1.2 $\alpha$-stable trajectories with rare infinite branchings
More precisely, for $\alpha\in(0,1)$, we define by $\nu_{\alpha}$ the normalized $\alpha$-stable Lévy measure on $(0,\infty)$, that is $\nu_{\alpha}(\mathrm{d}\mathfrak{z}):=\frac{\alpha}{\Gamma(1-\alpha)}\frac{\mathrm{d}\mathfrak{z}}{\mathfrak{z}^{\alpha+1}}$. We denote by $\tilde{\nu}_{\alpha}$ the natural extension of $\nu_{\alpha}$ to $\mathscr{Q}$ that consists in choosing the coordinate $x_{1}$ according to $\nu_{\alpha}$ and setting the other coordinates at $\infty$. More precisely, $\tilde{\nu}_{\alpha}(\mathrm{d}\bm{x}):=\nu_{\alpha}(\mathrm{d}x_{1})\otimes\delta_{\infty}(\mathrm{d}x_{2})\otimes\delta_{\infty}(\mathrm{d}x_{3})\otimes\ldots$ and, for $\beta\in(0,1)$, let $y^{\beta}:=(y^{\beta}_{n})_{n\geq 1}$ be defined by $y^{\beta}_{n}:=(\log n)^{\beta}$. We set $\Lambda^{ri}_{\alpha,\beta}(\mathrm{d}\bm{x})=\tilde{\nu}_{\alpha}(\mathrm{d}\bm{x})+\delta_{y^{\beta}}(\mathrm{d}\bm{x})$. In other words, each trajectory branches at rate $1$, at a branching time, displacements are given by the sequence $y^{\beta}$, and each trajectory follows a normalized $\alpha$-stable subordinator between branching times. Note that for any $\alpha,\beta\in(0,1)$, the conditions (3) and (5) are satisfied by $\Lambda^{ri}_{\alpha,\beta}$. This implies that the branching subordinator with characteristic couple $(0,\Lambda^{ri}_{\alpha,\beta})$ is well-defined, see Appendix A. Finally, note that the measure $\Lambda^{ri}_{\alpha,\beta}$ falls in the case (9).
@@ 1.1.3 $\alpha$-stable trajectories with many binary branchings
More precisely, for $\alpha\in(0,1)$ we let $\tilde{\nu}_{\alpha}$ be as in Section 1.1.2. For $\theta>1$, we let $\mu_{\theta}$ be the sigma-finite measure on $\mathbb{R}_{+}$ defined by $\mu_{\theta}(\mathrm{d}\mathfrak{z}):=e^{\mathfrak{z}^{\theta}}\mathds{1}_{\{\mathfrak{z}\geq 0\}}\mathrm{d}\mathfrak{z}$. We denote by $\tilde{\mu}_{\theta}$ the extension of $\mu_{\theta}$ to $\mathscr{Q}$ that consists in choosing the coordinate $x_{1}$ as 0, the coordinate $x_{2}$ according to $\mu_{\theta}$, and setting the other coordinates at $\infty$. More precisely, $\tilde{\mu}_{\theta}(\mathrm{d}\bm{x}):=\delta_{0}(\mathrm{d}x_{1})\otimes\mu_{\theta}(\mathrm{d}x_{2})\otimes\delta_{\infty}(\mathrm{d}x_{3})\otimes\delta_{\infty}(\mathrm{d}x_{4})\otimes\ldots$ and we set $\Lambda^{mb}_{\alpha,\theta}(\mathrm{d}\bm{x})=\tilde{\nu}_{\alpha}(\mathrm{d}\bm{x})+\tilde{\mu}_{\theta}(\mathrm{d}\bm{x})$. Note that for any $\alpha\in(0,1)$ and $\theta>1$, the conditions (3) and (5) are satisfied by $\Lambda^{mb}_{\alpha,\theta}$. This implies that the branching subordinator with characteristic couple $(0,\Lambda^{mb}_{\alpha,\theta})$ is well-defined, see Appendix A.
The measure $\Lambda^{mb}_{\alpha,\theta}$ clearly falls in the case (9) so this example illustrates that, for branching subordinators, one can very well have only binary branchings and yet be in the case (9), which would not be possible for branching random walks. This shows that the setting of branching subordinators is somehow more adapted than the setting of branching random walks to generate natural examples in the case (9) which, as we will see in Section 1.3, allows for a larger variety of asymptotic behaviors, compared with the classical case (6).
@@ 1.2 Some more definitions
The function $\kappa_{a}(\cdot)$, $a>0$, has a natural interpretation. For any $\bm{x}=(x_{i})_{i\geq 1}\in\mathscr{Q}$, we set $T_{a}:\bm{x}\in\mathscr{Q}\mapsto\bm{x}^{a}\in\mathscr{Q}$ to be the $a$-truncated operator on $\mathscr{Q}$, where $\bm{x}^{a}$ is defined to be the sequence $\bm{x}$ capped at $a$ such that $x_{1}$ remains unchanged, that is $(\bm{x}^{a})_{1}:=x_{1}$ and for any $i\geq 2$, $(\bm{x}^{a})_{i}:=x_{i}\mathds{1}_{\{x_{i}<a\}}+\infty\mathds{1}_{\{x_{i}\geq a\}}$. Let $\Lambda^{a}(\mathrm{d}\bm{x}):=\mathds{1}_{\{\bm{x}\neq(0,\emptyset)\}}(T_{a}\Lambda)(\mathrm{d}\bm{x})$, where $T_{a}\Lambda$ denotes the image measure of $\Lambda$ by $T_{a}$. Note that $\Lambda^{a}(\mathscr{Q}^{\star})=\Lambda(\{\bm{x}\in\mathscr{Q};x_{2}<a\})$ so $\lim_{a\to\infty}\Lambda^{a}(\mathscr{Q}^{\star})=\Lambda(\mathscr{Q}^{\star})$ which is positive by (3). Therefore, there exists $a_{0}(\Lambda)\geq 0$ such that $\Lambda^{a}(\mathscr{Q}^{\star})>0$ for all $a>a_{0}(\Lambda)$. Note that if $\Lambda$ satisfies (3) and (4), then $\Lambda^{a}$ satisfies the conditions (3) and (6) for all $a>a_{0}(\Lambda)$. For a branching subordinator $Y$ with characteristics $(d,\Lambda)$ and $a>a_{0}(\Lambda)$, we denote by $Y^{a}$ the sub-system of $Y$ obtained as follows: at each branching event (represented by an $\bm{x}=(x_{i})_{i\geq 1}\in\mathscr{Q}$ with $x_{2}<\infty$), the particle born with displacement $x_{1}$ is kept and for $i\geq 2$, the particles born with displacement $x_{i}\geq a$ are removed from the system, along with their future lineage. Then, as shown in Appendix A, $Y^{a}$ is a branching subordinator with characteristics $(d,\Lambda^{a})$ and the corresponding Laplace exponent is $\kappa_{a}(\cdot)$, that is, for any $t,\lambda\geq 0$
$\mathbb{E}\Big[\sum_{u\in\mathscr{N}^{a}(t)}e^{-\lambda Y^{a}_{u}(t)}\Big]=e^{-t\kappa_{a}(\lambda)}\in(0,\infty),$ (13)
where, similarly as for $Y$, $\mathscr{N}^{a}(t)$ stands for the set of particles in $Y^{a}$ alive at time $t$ and for any particle $u\in\mathscr{N}^{a}(t)$ and $s\leq t$, $Y^{a}_{u}(s)$ denotes the position of $u$ at time $s$, if already born at time $s$. Otherwise, $Y^{a}_{u}(s)$ stands for the position of the most recent ancestor of $u$ alive at time $s$. We refer to Appendix A for a joint construction of $Y$ and $Y^{a}$ where $Y^{a}$ is introduced first in Definition A.1 (and a justification of (13) is provided), $Y$ is subsequently constructed in Definition A.4, and a slightly different version of the above restriction property is stated in Lemma A.3. Also note that each canonical trajectory associated with $Y^{a}$ is a subordinator with Laplace exponent $\phi$, see below (3) for a description of canonical trajectories and Definition A.2 for a rigorous definition.
@@ 1.3 Main results
The main purpose of the present paper is to study the behavior of the leftmost particle of branching subordinators. This behavior is of particular interest to us because it turns out that, under our assumptions, for a branching subordinator $Y$ with characteristics $(d,\Lambda)$, the growth of $t\mapsto\underline{Y}(t)=\min_{u\in\mathscr{N}(t)}Y_{u}(t)$ can possibly display interesting sub-linear behaviors.
1.3.1 Divergence of the leftmost particle
Introduce $\underline{Y}(\infty):=\lim_{t\to\infty}\underline{Y}(t)$. Note that $\underline{Y}(\infty)$ is well defined since the function $t\mapsto\underline{Y}(t)$ is non-decreasing. We start by giving a criterion in terms of the truncated Laplace exponent to determine whether $\underline{Y}(\infty)$ is finite or infinite. For this, we first note from (4) and dominated convergence that, for any $0<a_{1}<a_{2}$, $M_{a_{2}}(\lambda)-M_{a_{1}}(\lambda)$ converges to 0 as $\lambda$ goes to infinity, so the quantity
@@ 1.3.1 Divergence of the leftmost particle
is independent of the choice of $a>0$. We note that $\mathscr{L}(d,\Lambda)$ can be simply expressed in terms of the characteristic couple $(d,\Lambda)$ of $Y$. Indeed, for any $\bm{x}\in\mathscr{Q}$ we set $N(\bm{x}):=\sharp\{n\geq 1;\ x_{n}=0\}$ and note that (4) implies in particular that $\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x})<\infty$. It is easy to see from (10), monotone convergence and dominated convergence that
$\mathscr{L}(d,\Lambda)=\infty\mathds{1}_{\{d>0\}}+\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})=0\})-\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x}),$ (15)
where $\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})=0\})\in[0,\infty]$.
@@
If the assumption (6) holds true for some $\theta\geq 0$ (we recall that in this case the Laplace exponent $\kappa(\cdot)$ introduced in (7) is well-defined on $\theta,\infty)$), then we have
$\mathscr{L}(d,\Lambda)=\sup_{\lambda\geq\theta}\kappa(\lambda)=\lim_{\lambda\rightarrow\infty}\kappa(\lambda)=\infty\mathds{1}_{\{d>0\}}+\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})=0\})-\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x}).$ (16)
Our criterion for finiteness or infiniteness of $\underline{Y}(\infty)$ is as follows.
@@
The analogue of the event $\{\underline{Y}(\infty)<\infty\}$ is referred to as the phenomenon of explosion in [1] and [3]. This terminology comes from the interpretation of the process as an age-dependent branching processes. Focused on branching random walks with non-negative displacements, authors of [1] provide conditions for explosion to happen. They also prove a result about the behavior of the leftmost particle in the case of non-explosion, that is when $\{\underline{Y}(\infty)=\infty\}$, see the end of section 1.3.2 of the present article. In our case, in view of Proposition 1.3, it is sufficient to assume $\mathscr{L}(d,\Lambda)>0$ to ensure that explosion does not occur.
1.3.2 Classical results and linear behavior
A natural question is that of the rate of growth of $\underline{Y}(t)$. In any case, under (3) and (4), we have, $\mathbb{P}$-almost surely
@@ 1.3.2 Classical results and linear behavior
As for branching random walks (see for example the famous works of Hammersley [18], Kingman [20] and Biggins [11]), if (6) holds for some $\theta\geq 0$ and if we further require that $\infty<\kappa(\theta_{0})<0$ for some $\theta_{0}>0$ (allowing $\kappa(0)=$-$\infty$), then the limit exists and is nonrandom, more precisely, we have $\mathbb{P}$-almost surely
$\lim_{t\to\infty}\frac{\underline{Y}(t)}{t}=\sup_{\lambda\geq\theta_{0}}\frac{\kappa(\lambda)}{\lambda}=\sup_{\lambda>0}\frac{\kappa(\lambda)}{\lambda}\in[0,\infty).$ (18)
Therefore, in the latter case, (18) provides more information about the behaviors described in Proposition 1.3: The growth of $t\mapsto\underline{Y}(t)$ is linear if $\mathscr{L}(d,\Lambda)>0$. The growth is sub-linear if $\mathscr{L}(d,\Lambda)=0$, even when $\underline{Y}(\infty)=\infty$; the branching Poisson process in the case $r=\rho$ provides such an example, see Section 3.2.1 for details. Finally, recall that $\underline{Y}(\infty)\in[0,\infty)$ if $\mathscr{L}(d,\Lambda)<0$.
Note that, by (17), the classical linear growth is observed when $d>0$, even when the assumptions ensuring (18) do not hold, for example in the case (9). We discuss the linear behavior and provide a brief justification of (17) and (18) in Appendix B. The justification of (18) consists in transposing to our setting the classical result obtained for branching random walks in the above mentioned references.
1.3.3 Relaxing the exponential moments assumption, sub-linear behavior
@@
We define a measure $\mu$ on $0,\infty)$ by
$\forall A\in\mathscr{B}(\mathbb{R}_{+}),\ \mu(A):=\int_{\mathscr{Q}}\left(\sum_{k=2}^{\infty}\mathds{1}_{\{x_{k}\in A\}}\right)\Lambda(\mathrm{d}\bm{x}).$ (20)
The measure $\mu$ can be understood as a measure-valued average of the contributions of branchings. We note that, for $a>0$, $M_{a}(\lambda)=\int_{[0,a)}e^{-\lambda\mathfrak{z}}\mu(\mathrm{d}\mathfrak{z})$. If we had $\mu([0,a))$ of order $e^{ca}$, then (6) would be satisfied for any $\theta>c$. Since we are mostly interested in the case (9), we assume an higher rate of growth for $\mu([0,a))$, but also some regularity. More precisely, our second assumption is the following.
@@
$M_{2,a}(\lambda):=\int_{\mathscr{Q}}\left(\sum_{k=2}^{\infty}e^{-\lambda x_{k}}\mathds{1}_{\{x_{k}<a\}}\right)^{2}\Lambda(\mathrm{d}\bm{x}).$ (22)
@@
The variance control from Assumption 3 will allow, in Section 4.2, to show that some functionals involved in estimating the deviation probability for the leftmost particle concentrate around their means, which are themselves expressed in terms of quantities of the form $M_{a}(ca^{\sigma})$ (for some $c$ and $a$) whose behavior can be inferred from Assumption 2. Without Assumption 3, it might be difficult to rigorously relate the behavior of the system to the quantities of the form $M_{a}(ca^{\sigma})$.
Remark 1.4 ($p$-ary branchings).
By Jensen inequality we have, for $p\geq 2$, $(\sum_{k=2}^{p}e^{-\lambda x_{k}}\mathds{1}_{\{x_{k}<a\}})^{2}\leq(p-1)\sum_{k=2}^{p}e^{-2\lambda x_{k}}\mathds{1}_{\{x_{k}<a\}}$. Therefore, if there is $p\geq 2$ such that the system has at most $p$-ary branchings, that is, if $\Lambda(\{\bm{x}\in\mathscr{Q};x_{p+1}<\infty\})=0$, then we have $M_{2,a}(\lambda)\leq(p-1)M_{a}(2\lambda)$ for any $a,\lambda\geq 0$ (if the system has only binary branchings then we even have $M_{2,a}(\lambda)=M_{a}(2\lambda)$). This will allow to conveniently check that Assumption 3 is satisfied in some examples involving binary or $p$-ary branchings (see Section 2.2 and Remark 2.2).
Our last assumption says that a given particle does not jump and give birth at the same time (even though its children can be born with positive displacement), so that jumps and branching events are independent.
Assumption 4.
We have $\Lambda(\{\bm{x}\in\mathscr{Q};0<x_{1}\leq x_{2}<\infty\})=0$.
The independence, provided by Assumption 4, of jumps and branching events allows for slow particles to not be penalized by the fact that they are slow, so that their branching structure is distributed as the branching structure of a generic particle.
@@
Let $\Lambda$ be a measure on $\mathscr{Q}$ satisfying (3)-(5) and Assumption 1 for some $\alpha\in(0,1)$, Assumption 2 for some $\sigma>0$, and Assumptions 3 and 4. Then the branching subordinator with characteristic couple $(0,\Lambda)$ satisfies (19) with $\gamma:=\frac{1}{\alpha+(1-\alpha)(1+\sigma)}$.
@@
We will see in Remark 2.2 that the result of Corollary 1.7 remains true with the same value of $\gamma$ when the binary branchings from the example of Section 1.1.3 are replaced by $p$-ary branchings.
The fact that the growth of $t\mapsto\underline{Y}(t)$ is slower than linear under our assumptions can informally be explained as follows: in cases where $\lim_{a\to\infty}M_{a}(\lambda)=\infty$ for any $\lambda\geq 0$, each particle in $Y$ is somehow allowed to give progeny to a very large number of children (see the example from Section 1.1.2 for instance) or/and to create newborn children very often (see the example from Section 1.1.3 for instance), producing a population of particles sufficiently large to maintain $\underline{Y}(t)$ below $\mathfrak{c}t$ for any $\mathfrak{c}>0$. Besides, one can notice that the rate of growth of $t\mapsto\underline{Y}(t)$ depends on both $\alpha$ (see Assumption 1) and $\sigma$ (see Assumption 2). The parameter $\alpha$ acts on the intensity of the (infinitely many) small jumps of each particle while $\sigma$ acts on the distribution of the positions of the children of each particle and the combination of both is largely responsible for this rate of growth.
The idea of the proof of Theorem 1.5 is as follows. Proving the lower bound relies on estimating the average contribution of slow particles. Proving the upper bound is more involved and relies on understanding the mechanism by which slow particles are generated and to show that, in some sense, the structure of branchings allows for the set of slow particles to sustain itself over time. In other words we produce events of large probability on which slow trajectories produce sufficiently enough slow particles. This requires to prove a large deviation estimate that gives a lower bound on the probability for a particle to be slow, and to show that there is some regularity on the distribution of particles that are born from the slow trajectories.
Properties of branching processes are often established through martingale technics. In the light of this, it seems that another approach might be possible to prove the upper bound in Theorem 1.5. While the martingale $(W_{\lambda}(t))_{t\geq 0}$ from (8) is not well-defined under the assumptions of Theorem 1.5, one can define an analogue of that martingale for the truncated process $Y^{a}$, where the truncation $a$ increases with time. Then, one might be able to proceed as in [13] (using a spinal decomposition argument) to determine suitable conditions for that martingale to have a positive limit. Then, if one can show that the contribution to that martingale of particles on the right of position $ct^{\gamma}$ (for some constant $c>0$) converges to zero almost surely, one can deduce the almost sure existence of particles on the left of $ct^{\gamma}$ at all large times $t$, yielding the result. In comparison, the approach that we used to prove Theorem 1.5 seems more direct, probably more concise, and provides more insights on the mechanism by which slow particles arise and give birth to new slow particles before speeding up.
Although the article [1] is mainly devoted to explosion, their Theorem 6.1 is dedicated to the case of non-explosion. Considering a branching random walk $X=(X(n))_{n\geq 1}$ with non-negative and i.i.d displacements, authors of [1] proved a rather precise result: assuming there is non-explosion for the random walk $X$, together with a heavy tail assumption on the offspring distribution, they show that, conditionally on the non-extinction of the system, the renormalized position of the leftmost particle $\underline{X}(n)/x_{n}$ converges to $1$ almost surely, where $(x_{n})_{n\geq 1}$ is a deterministic semi-explicit sequence of positive numbers. We note that authors of [1] considered i.i.d displacements (that assumption plays an important role in the definition of their renormalizing sequence $(x_{n})_{n\geq 1}$) which is a setting completely different from the cases we consider.
1.3.4 Organization of the paper